OpenStudy (anonymous):

Solve: 4(x + 1) = 4x + 4 No Solution 0 Any Real Number 1

OpenStudy (anonymous):

any real number

OpenStudy (anonymous):

you sure why?

OpenStudy (anonymous):

coz if u put any real no. into the equation then both sides will be equal

OpenStudy (anonymous):

I thought its no solution or 1

OpenStudy (anonymous):

4(1+1)=4(1)+4 4(-3+1)=4(-3)+4 8=8 -8=-8 since we r fnding the value of x it is a set of real no.s becoz ny number will make the statement true

OpenStudy (anonymous):

ahh I see thankx